If the forecast for September is based on the data series below and the forecast for September is 19, then which forecasting method was used?

January February March April May June July August
18 21 20 22 21 20 21 17

Moving average (n = 2)

Moving average (n = 3)

Averaging

Last value
Answer #1

Question: If the forecast for September is based on the data series below and the forecast for September is 19, then which forecasting method was used?

Answer: Moving Average (n = 2)

Explanation:

1. Moving Average is given by:

Moving Average = (A1 + A2 + A3 + ... + An) / n

Here for:

n = 2

September Forecast = (17 + 21) / 2 = 19

n = 3

September Forecast = (17 + 21 + 20) = 19.33

2. Averaging

September Forecast = (17 + 21 + 20 + 21 + 22 + 20 + 21 + 18) / 8 = 20

3. Last Value

Last Values is given by:

Forecast Ft = Actual At-1

Therefore:

September Forecast = August Actual = 17

Conclusion: Moving Average Forecast with n = 2 provides the September forecast = 19